OpenStudy (anonymous):

NEED HELP with question #8

OpenStudy (eyust707):

whats question 8?

hero (hero):

You do realize that the t^3 - t = theta, right?

OpenStudy (anonymous):

ya

hero (hero):

So you can actually set t^3 - t = 90 and solve for t. you already know that sin(90) = 1, so that means that if you find t, you'll find the t where x = 1

hero (hero):

This is a little tricky though. Solving for t won't be as simple as that

OpenStudy (anonymous):

no they plugged in values to one that was similar to this

OpenStudy (anonymous):

like 1,0 and 100 but that was when t was an exponent

OpenStudy (anonymous):

shld I skip this one

hero (hero):

For instance, when t is approximately 4.5, x = 1, theta = 90

hero (hero):

No, you shouldn't.

OpenStudy (anonymous):

well i want to

hero (hero):

I solved that one using a calculator because that's the only way I would know how to solve these

hero (hero):

but, lets do one that you would know how to do

hero (hero):

t^3 - t = theta when theta = 0

hero (hero):

Do that one

OpenStudy (anonymous):

I have to hand in this question it is evn

hero (hero):

rld, why are you doing even numbered questions. You're wasting both our time

OpenStudy (anonymous):

no that is what i need to hand in

OpenStudy (anonymous):

t=1

OpenStudy (anonymous):

or 0

hero (hero):

it's corresponding to angles instead of time. I know it's a natural progression to think that t always responds to time, but it doesn't

OpenStudy (anonymous):

oh right

hero (hero):

You have to finish it though t can equal 1 or zero either way theta = 0 so sin(0) = 0 cos(0) = 1 so x = 0, y = 1

OpenStudy (anonymous):

ya but what does t=

hero (hero):

t = -1, 0, or 1, remember? You calculated it

OpenStudy (anonymous):

i went on to the next one already

hero (hero):

You have to find t, x, and y

hero (hero):

not just t

OpenStudy (anonymous):

I wanna go to bed already

hero (hero):

lol

OpenStudy (eyust707):

lol

hero (hero):

Nobody is stopping you from going to bed

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

idk

OpenStudy (eyust707):

for these i just plug in t=0 and it gives me an x and it gives me a y then i do t=1 t=2 t=3...

OpenStudy (anonymous):

oh ok

hero (hero):

Yeah, I kind of went backwards with it

OpenStudy (anonymous):

culate it in degrees or radians?

OpenStudy (anonymous):

culate it in degrees or radians?

OpenStudy (eyust707):

well actually for these we can pick any t we want

hero (hero):

degrees

OpenStudy (eyust707):

so its easier to pick t's that are easy to solve

OpenStudy (anonymous):

oh man that is where i was going wrong

hero (hero):

I hate radians

OpenStudy (anonymous):

u hate e/t

OpenStudy (eyust707):

i would use t=-1 t=0 and t=1 as a start

OpenStudy (anonymous):

ok

hero (hero):

it doesn't say whether to use radians or degrees

hero (hero):

any input on that eyust

OpenStudy (eyust707):

well it doesnt really matter.

OpenStudy (eyust707):

so it wants a parameterization

OpenStudy (anonymous):

my book likes to use radians but i dont care i will use what i want to

hero (hero):

yeah, I don't think it matters

OpenStudy (eyust707):

and it tell you its a circle

OpenStudy (anonymous):

yaa

OpenStudy (eyust707):

so a circle is x^2 + y^2 = r^2

OpenStudy (eyust707):

we know x and we know y

OpenStudy (eyust707):

find r by pulgging in your t values

hero (hero):

actually, we still have to figure out x and y. We're inputting t

OpenStudy (anonymous):

huh we didn't do it like that

OpenStudy (eyust707):

(cost^3-t)^2 + (sint^3-t) ^2 = r^2

hero (hero):

Oh, lookie there

hero (hero):

You still have to find x and y

OpenStudy (eyust707):

r = 1

hero (hero):

making it look pretty doesn't change the fact that you still have to calculate the cosine and sine based on t that gives you x and y that you need to solve it

OpenStudy (eyust707):

a parameterization just says change the function of x and y to just t's

hero (hero):

Okay, that's good. My way is just the long way to the same goal. Your approach is superior, no doubt.

OpenStudy (eyust707):

and also tell us when t starts and stops

OpenStudy (anonymous):

ya be nice to eyust

hero (hero):

You assume that I was being mean to him. rld, you don't need to tell me what to do. I can handle myself

hero (hero):

I have a feeling I will get to bed before rld does

OpenStudy (anonymous):

R u ready to kill me @hero

hero (hero):

Now why would I want to do that?

OpenStudy (anonymous):

cuz i am annoying

OpenStudy (eyust707):

ok so @ t= 0 we get (1,0) but the graph for this is nuts!

OpenStudy (eyust707):

it ends up being a solid circle with radius 1 but here let me see if w.a will graph it

OpenStudy (anonymous):

I also got a crazy graph

OpenStudy (eyust707):

ask you professor/teacher about number 8 i honestly am not sure what they want there

OpenStudy (eyust707):

get some rest imma do the same =P

OpenStudy (eyust707):

later man

OpenStudy (anonymous):

what time is it by u

OpenStudy (anonymous):

thanks for ur help

OpenStudy (eyust707):

12 44 got class at 8